# Curious QNEIs from QNEC: New Bounds on Null Energy in Quantum Field Theory

### v1

@@ 3.2 A family of interacting QNEIs
where $V_{\perp}\equiv\int\text{d}^{d-2}y_{\perp}$ is the IR-divergent transverse volume. To be careful and clear in this null limit, (97) represents the divergent and finite terms in a $\epsilon_{B}$ expansion while holding $v-v_{B}$ fixed such that $\epsilon_{B}|v-v_{B}|>\epsilon_{\text{UV}}^{2}$, where $\epsilon_{\text{UV}}$ is an inverse UV cutoff. We have additionally allowed for a $(u,v)$ independent divergence with the UV cutoff, $\epsilon_{\text{UV}}$ stemming from a normalization of the $D_{n}D_{n}$ OPE; however, importantly, there are no further subleading divergences in $\epsilon_{\text{UV}}$. Subleading terms would be functionals of curvature invariants of the entanglement cut, [^15] The general structure of the vacuum EE of any region $\Sigma$ in any local QFT is Casini:2022rlv $S_{vac.}(\Sigma)=\sum C_{i}(\partial\Sigma)\epsilon_{\text{UV}}^{-\lambda_{i}}+S_{0}(\Sigma)$ (98) where $C_{i}(\partial\Sigma)$ have dimension $\lambda_{i}$ and are local and extensive functionals of the curvature invariants of the boundary $\partial\Sigma$: $C_{i}(\partial\Sigma)=\int_{\partial\Sigma}(\dots)$ (99) which vanish for our flat entanglement cuts. For holographic CFTs, the vacuum entropy of a strip appeared as an example in the original paper by Ryu and Takayanagi, Ryu:2006bv , however, to our knowledge, the universality of this result for general interacting CFTs in the null limit has not yet been emphasized in the literature.
@@
$\partial_{v}^{\scriptscriptstyle{(\uparrow)}}S_{\Omega}(B)=\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}\,V_{\perp}}{\epsilon_{B}(u)^{\frac{d-2}{2}}(v-v_{B})^{\frac{d}{2}}}+O(\epsilon_{B})~.$ (100)
@@
As we mentioned, (3.2) supposes that the only operator saturating the twist gap is the single-copy stress tensor, $\sum_{i}T_{vv}^{(i)}$. However, in free theories, this assumption is not true. For instance, there exist an infinite tower of single-copy higher-spin operators, schematically,
@@
for all even $s$ which can alter the expression for $\Delta S$. In this case, $\Delta S\neq\Delta K$ and we need to instead use the monotonicity of the relative entropy (as in Section 2.2) to isolate an inequality involving only the stress tensor. Moreover there are also double-copy higher-spin operators of the form
@@
$\partial_{v}S(A)\leq\int\text{d}^{d-2}y_{\perp}$ $\left[2\pi\int_{v_{B}}^{v}\text{d}v^{\prime}\mathcal{G}\left(\frac{v^{\prime}-v_{B}}{v-v_{B}}\right)\langle T_{vv}(u,v^{\prime},y_{\perp})\rangle+\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}}{\epsilon_{B}(u)^{\frac{d-2}{2}}(v-v_{B})^{\frac{d}{2}}}\right]$
@@
$\partial_{v}S(A)\geq\int\text{d}^{d-2}y_{\perp}$ $\left[-2\pi\int_{v}^{v_{C}}\text{d}v^{\prime}\mathcal{G}\left(\frac{v_{C}-v^{\prime}}{v_{C}-v}\right)\langle T_{vv}(u,v^{\prime},y_{\perp})\rangle-\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}}{\epsilon_{C}(u)^{\frac{d-2}{2}}(v_{C}-v)^{\frac{d}{2}}}\right]$
@@
We can now follow similar steps to those in Section 2.2. That is, we return to the integral (64) and we will suppose $g(u,v)$ to be such that $\text{sgn}(v)\partial_{v}g\geq 0$ and we will define functions
$\zeta(u,v)=\begin{cases}v_{B}(u,v),&v<0\\ v_{C}(u,v),&v\geq 0\\ \end{cases}~,\qquad\epsilon_{u}(u,v)=\begin{cases}\epsilon_{B}(u,v),&v<0\\ \epsilon_{C}(u,v),&v\geq 0\\ \end{cases}~.$ (105)
More generally, we can take $\zeta(u,v)$ and $\epsilon_{u}(u,v)$ to be any functions such that
$\text{sgn}(v)\text{sgn}(\zeta(u,v)-v)\geq 0~,\qquad-\text{sgn}(v)\text{sgn}(\epsilon_{u}(u,v))\geq 0~.$ (106)
Next we invert the integration domain to arrive at
$\int\text{d}^{d-2}y_{\perp}\int\text{d}u\text{d}v$ $\mathcal{M}(u,v)\,\langle T_{vv}\rangle\geq$
$-\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}\,V_{\perp}}{2\pi}\int\text{d}u\text{d}v\,\left(\frac{\partial_{v}g(u,v)}{\epsilon_{u}(u,v)^{\frac{d-2}{2}}(v-\zeta(u,v))^{\frac{d}{2}}}+O(\epsilon_{u})\right)~,$ (107)
@@
$\mathcal{M}(u,v):=$ $g(u,v)-\int_{\zeta^{-1}(u,v)}^{v}\,\text{d}v^{\prime}(\partial_{v^{\prime}}g(u,v^{\prime}))\mathcal{G}\left(\frac{v-\zeta(u,v^{\prime})}{v^{\prime}-\zeta(u,v^{\prime})}\right)$
$=$ $-\frac{1}{2}\int_{\zeta^{-1}(u,v)}^{v}\text{d}v^{\prime}\,g(u,v)\,G^{\prime\prime}\left(\frac{v-\zeta(u,v^{\prime})}{v^{\prime}-\zeta(u,v^{\prime})}\right)\frac{\text{d}}{\text{d}v^{\prime}}\left(\frac{v-\zeta(u,v^{\prime})}{v^{\prime}-\zeta(u,v^{\prime})}\right)^{2}~.$ (108)
We see that the function $\mathcal{M}$ takes a form very similar to the function appearing in our 2d QNEIs (59) with the exception of its defining integral being weighted by the second derivative of the kernel appearing in the modular Hamiltonian associated to a null strip, (90). Note that for free field theories $G^{\prime\prime}(s)=-2$ and so $\mathcal{M}$ (at a fixed $u$) coincides with $m$. More generally though, because of the unspecified theory-specific nature of this modular kernel, it is hard to establish general properties of $\mathcal{M}$. However we do point out that when $G$ is concave, i.e. $G^{\prime\prime}(s)\leq 0$, as was conjectured in Bousso:2014uxa , then we can easily establish that $\mathcal{M}$ obeys bounds analogous to (60), namely
$g(u,v)\leq\mathcal{M}(u,v)\leq g(u,\zeta^{-1}(v))~.$ (109)
@@
$\int\text{d}^{d-2}y_{\perp}$ $\int\text{d}u\,\text{d}v\,\mathcal{M}(u,v)\langle T_{vv}(u,v,\vec{y}_{\perp})\rangle\geq$
$-\frac{\beta_{d}c_{\scriptscriptstyle\mathrm{T}}V_{\perp}}{2\pi}\int\text{d}u\,\text{d}v\left(\frac{\partial_{v}\mathcal{M}(u,\zeta)\Theta(\partial_{v}\zeta-1)+\partial_{v}\mathcal{M}(u,v)\Theta(1-\partial_{v}\zeta)}{\epsilon_{u}^{\frac{d-2}{2}}(v-\zeta)^{\frac{d}{2}}}+O(\epsilon_{u})\right)$ (110)
where for brevity of notation we have left the $(u,v)$ dependence of $\zeta$ implicit.
Again, much like our simplified 2d QEI, (61), (3.2) is now expressed in terms of a single smearing function, $\mathcal{M}$, and two functions, $\zeta$ and $\epsilon_{u}$ parameterizing choices of nearly null intervals. For a given $G$, which is fixed by the theory in question, it is unclear to us whether any generic smearing function $\mathcal{M}$ can be realized by a choice of $g$ appearing in (3.2). If so, then (61) and (3.2) would represent a very tantalizing result: a universal family of QNEIs for interacting CFTs which, up to overall constant, are independent of the particular details of a CFT.
@@ 4 Discussion
In this paper, we have derived new quantum null energy inequalities, i.e. lower bounds on $\langle T_{vv}\rangle_{\psi}$, the null-null component of the stress tensor, integrated over a spacetime region in interacting quantum field theories. While our study of QNEIs is partially motivated by their potential applications to semi-classical gravity (e.g. as input in a semi-classical singularity theorme) such lower bounds on null energy are, of course, of intrinsic interest in the study of interacting QFTs. In particular the foundation of our QNEIs are deep connections between energy and entropy: in our derivation we utilize QNEC, strong subadditivity, and the vacuum modular Hamiltonians of null strips. More specifically, given a positive smearing function, $g$, we bound the stress tensor integrated against a smooth kernel, determined by the properties of the modular Hamiltonian of a null strip, in terms of a state-independent functional of $g$. In the case of $d>2$ dimensions, our bounds are the first such state-independent QEIs integrated over a finite null segment in a broad class of interacting QFTs, namely CFTs.
@@
where a precise form of $m$ can be found in (59). In contrast to QNEC, these lower bounds are state-independent, because the right hand side does not depend on the state $\psi$. And, in contrast to ANEC, these bound are localised, in the sense that we can choose $g$ such that we are integrating $\langle T_{vv}\rangle$ against a bump function on the left hand side.
We have an infinite family of QNEIs, one for each choice of $\zeta$, which is both powerful but also increases the complexity of the inequality. The choice of $\zeta$ affects both the definition of $m$, the correction to the smearing function, and the value of the lower bound and arises from a choice of null intervals appearing in intermediate stages of our derivation. However, ultimately $\zeta$ is an invertible function obeying (48), but otherwise arbitrary. For a given $\langle T_{vv}\rangle$, there does exist a $\zeta$ that makes this intermediate stage inequality as tight as possible, but, for general $\langle T_{vv}\rangle$, there is not an analytic solution for this $\zeta$ (see Sec. 2.2.4). We explicitly determined the QNEI for a few example choices of $\zeta$. It would be interesting to see whether the general form of our QNEIs can be simplified.
@@
$\begin{split}\int\text{d}^{d-2}y_{\perp}&\int\text{d}u\text{d}v\mathcal{M}(u,v)\langle T_{vv}\rangle\\ &\geq-\frac{\beta\,c_{\scriptscriptstyle\mathrm{T}}}{2\pi}\int\text{d}^{d-2}y_{\perp}\int\text{d}u\text{d}v\left(\frac{\partial_{v}g(u,v)}{\epsilon_{u}(u,v)^{\frac{d-2}{2}}(v-\zeta(u,v))^{\frac{d}{2}}}+O(\epsilon_{u})\right)~,\end{split}$ (112)
@@
where $\mathcal{M}$ is defined in (3.2). As for $d=2$, this inequality is state-independent, and the integral over $\langle T_{vv}\rangle$ is localised (in the $(u,v)$ plane). Unlike for $d=2$, we have to integrate $\langle T_{vv}\rangle$ over the second null direction $u$ to get a non-trivial bound and over the transverse directions $y_{\perp}$ to get a universal result. Additionally this bound depends on a second function, $\epsilon_{u}$, which parametrises how close we take our strip regions to being null in an intermediate step of the derivation (see Sec. (3.2)), and we have to work perturbatively in this function, to make use of the universal form of $\Delta K$ for null strips. To our knowledge, these are the first QEIs, integrated over a finite null region, proven for higher-dimensional interacting QFTs.
@@
While we view this work as an important first step in the study of QNEIs in interacting theories there is still much more to explore:
@@
  • Our QNEIs in higher dimensions are localised in the $(u,v)$ plane, but not in the transverse space, so they are not truly localised bounds on the stress tensor. Additionally this integration strictly trivialises the right-hand side of our bounds (albeit in a physically uniteresting way). It would obviously be interesting and useful to try to leverage the techniques we’ve established here to find bounds that are also localized in this transverse direction. At a technical level, it was necessary to have a flat entanglement cut in the transverse directions to make use of the results for $\Delta K$ for a null strip given in Bousso:2014uxa . Having the expression for $\Delta K$ for null strips with “wiggly edges” would allow us to utilize local variations and the QNEC in its local form, (69). However this expression for the modular Hamiltonian is not known outside of free theories. The modular Hamiltonians of infinite null half-sheets with wiggly edges do admit closed form expressions Faulkner:2016mzt ; Casini:2017roe however we have so far not been able to leverage this expression to derive any new interesting bounds. Regardless, it is worth revisiting and reviewing this approach more carefully in the future.
@@
    It is also useful to look for alternative approaches for deriving a truly local QNEI. One potential approach is trying to realize the stress tensor in the OPE of two operators. This is essentially a more sophisticated form of the free field theory statement that $T_{vv}$ is a square of two operators. One obvious hang-up to this approach is that operator spectra in CFTs are not universal and so there is potential ambiguity in which OPE we try to realize stress tensor. To address this we can once again utilize entanglement based techniques. Namely defect CFTs living along twist defects defining replica path-integrals admit a universal set of displacement operators with the stress tensor appearing in their OPE channel Balakrishnan:2019gxl . One might speculatively hope that a reflection symmetric OPE of such displacement operators could provide a universal and local QEI in interacting CFTs.
@@
  • We have worked in Minkowski spacetime, both for simplicity and because our QNEIs are only proven for theories and background geometries for which our inputs are proven, QNEC in particular. The proof of QNEC given in Balakrishnan:2017bjg is valid for all relativistic QFTs in Minkowski spacetime. However, there are also proofs of QNEC on curved backgrounds, for a smaller set of QFTs, such as in Bousso:2014sda ; Fu:2017evt . It would advantageous, particularly with the application to semi-classical gravity in mind, to generalise our work to curved backgrounds. One obstacle to overcome is the transverse integration of the previous bullet point. Because the integration region of our bound in $d>2$ dimensions could detect global structures of a background, we might incur large curvature corrections or even topological obstructions to our bound. A more local expression of the form mentioned in the previous bullet would allow us to work perturbatively when smearing over regions smaller than the typical curvature scale.
@@
  • Our derivations take QNEC as one of their inputs. In the same way that, while ANEC can be proven by integrating QNEC along a complete null line, it can also be proven directly using causality Hartman:2016lgu and monotonicity of relative entropy Faulkner:2016mzt , so we expect that our QNEIs can be proven without taking QNEC as input, perhaps using the causality and monotonicity of relative entropy.
@@
  • It would be interesting to extend our work to QFTs coupled to gravity. One intriguing possibility is that gravity itself provides a natural regulator of negative energy densities. This is essentially the core idea of the SNEC (4) mentioned in the introduction of this paper. However, as mentioned there, collective evidence for or even proving such QEIs is a hard task because they require coupling to gravity as input. Existing proofs of QNEC, such as Balakrishnan:2017bjg , assume a fixed background geometry, but the original statement of QNEC came as a corrollorary of the QFC Bousso:2015mna , and the motivation of that paper was to study QFTs coupled to gravity. It would be interesting to see whether QEIs for QFTs coupled to gravity can be proven directly from the QFC. We are actively working in this direction and will report results in a future publication gravRolphBounds .

### v2

@@ 3.2 A family of interacting QNEIs
where $V_{\perp}\equiv\int\text{d}^{d-2}y_{\perp}$ is the IR-divergent transverse volume. To be careful and clear in this null limit, (99) represents the divergent and finite terms in a $\epsilon_{B}$ expansion while holding $v-v_{B}$ fixed such that $\epsilon_{B}|v-v_{B}|>\epsilon_{\text{UV}}^{2}$, where $\epsilon_{\text{UV}}$ is an inverse UV cutoff. We have additionally allowed for a $(u,v)$ independent divergence with the UV cutoff, $\epsilon_{\text{UV}}$ stemming from a normalization of the $D_{n}D_{n}$ OPE; however, importantly, there are no further subleading divergences in $\epsilon_{\text{UV}}$. Subleading terms would be functionals of curvature invariants of the entanglement cut, [^14] The general structure of the vacuum EE of any region $\Sigma$ in any local QFT is Casini and Huerta (2023) $S_{vac.}(\Sigma)=\sum C_{i}(\partial\Sigma)\epsilon_{\text{UV}}^{-\lambda_{i}}+S_{0}(\Sigma)$ (100) where $C_{i}(\partial\Sigma)$ have dimension $\lambda_{i}$ and are local and extensive functionals of the curvature invariants of the boundary $\partial\Sigma$: $C_{i}(\partial\Sigma)=\int_{\partial\Sigma}(\dots)$ (101) which vanish for our flat entanglement cuts. For holographic CFTs, the vacuum entropy of a strip appeared as an example in the original paper by Ryu and Takayanagi, Ryu and Takayanagi (2006), however, to our knowledge, the universality of this result for general interacting CFTs in the null limit has not yet been emphasized in the literature.
@@
$\partial_{v}^{\scriptscriptstyle{(\uparrow)}}S_{\Omega}(B)=\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}\,V_{\perp}}{\epsilon_{B}^{\frac{d-2}{2}}(v-v_{B})^{\frac{d}{2}}}+O(\epsilon_{B})~.$ (102)
@@
As we mentioned, (3.2) supposes that the only operator saturating the twist gap is the single-copy stress tensor, $\sum_{i}T_{vv}^{(i)}$. However, in free theories, this assumption is not true. For instance, there exists an infinite tower of single-copy higher-spin operators, schematically,
@@
for all even $s$ which can alter the expression for $\Delta S$. In this case, $\Delta S\neq\Delta K$ and we need to instead use the monotonicity of the relative entropy (as in Section 2.2) to isolate an inequality involving only the stress tensor. Moreover, there are also double-copy higher-spin operators of the form
@@
$\partial_{v}S(A)\leq\int\text{d}^{d-2}y_{\perp}$ $\left[2\pi\int_{v_{B}}^{v}\text{d}v^{\prime}\mathcal{G}\left(\frac{v^{\prime}-v_{B}}{v-v_{B}}\right)\langle T_{vv}(u_{0},v^{\prime},y_{\perp})\rangle+\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}}{\epsilon_{B}(v)^{\frac{d-2}{2}}(v-v_{B})^{\frac{d}{2}}}\right]$
@@
$\partial_{v}S(A)\geq\int\text{d}^{d-2}y_{\perp}$ $\left[-2\pi\int_{v}^{v_{C}}\text{d}v^{\prime}\mathcal{G}\left(\frac{v_{C}-v^{\prime}}{v_{C}-v}\right)\langle T_{vv}(u_{0},v^{\prime},y_{\perp})\rangle-\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}}{\epsilon_{C}(v)^{\frac{d-2}{2}}(v_{C}-v)^{\frac{d}{2}}}\right]$
@@
We can now follow similar steps to those in Section 2.2. Again, we restrict $g(v)$ to be such that $\text{sgn}(v)\partial_{v}g\geq 0$, and we define functions
$\zeta(v)=\begin{cases}v_{B}(v),&v<0\\ v_{C}(v),&v\geq 0\\ \end{cases}~,\qquad\epsilon_{u}(v)=\begin{cases}\epsilon_{B}(v),&v<0\\ \epsilon_{C}(v),&v\geq 0\\ \end{cases}~.$ (107)
More generally, we can take $\zeta(v)$ and $\epsilon_{u}(v)$ to be any functions such that
$\text{sgn}(v)\text{sgn}(\zeta(v)-v)\geq 0~,\qquad-\text{sgn}(v)\text{sgn}(\epsilon_{u}(v))\geq 0~.$ (108)
Next, we invert the integration domain to arrive at
$\int\text{d}^{d-2}y_{\perp}\int\text{d}v$ $\mathcal{M}(v)\,\langle T_{vv}\rangle\geq$
$-\frac{\beta_{d}\,c_{\scriptscriptstyle\mathrm{T}}\,V_{\perp}}{2\pi}\int\text{d}v\,\left(\frac{\partial_{v}g(v)}{\epsilon_{u}(v)^{\frac{d-2}{2}}(v-\zeta(v))^{\frac{d}{2}}}+O(\epsilon_{u})\right)~,$ (109)
@@
$\mathcal{M}(v):=$ $g(v)-\int_{\zeta^{-1}(v)}^{v}\,\text{d}v^{\prime}(\partial_{v^{\prime}}g(v^{\prime}))\mathcal{G}\left(\frac{v-\zeta(v^{\prime})}{v^{\prime}-\zeta(v^{\prime})}\right)$
$=$ $-\frac{1}{2}\int_{\zeta^{-1}(v)}^{v}\text{d}v^{\prime}\,g(v^{\prime})\,G^{\prime\prime}\left(\frac{v-\zeta(v^{\prime})}{v^{\prime}-\zeta(v^{\prime})}\right)\frac{\text{d}}{\text{d}v^{\prime}}\left(\frac{v-\zeta(v^{\prime})}{v^{\prime}-\zeta(v^{\prime})}\right)^{2}~.$ (110)
We see that the function $\mathcal{M}$ takes a form very similar to the function appearing in our 2d QNEIs (59) with the exception of its defining integral being weighted by the second derivative of the kernel appearing in the modular Hamiltonian associated to a null strip, (92). Note that for free field theories $G^{\prime\prime}(s)=-2$, in which case $\mathcal{M}$ coincides with $m$. More generally though, because of the theory-dependent nature of this modular kernel, it is hard to establish general properties of $\mathcal{M}$. However we do point out that when $G$ is concave, i.e. $G^{\prime\prime}(s)\leq 0$, as was conjectured in Bousso et al. (2015), then we can easily establish that $\mathcal{M}$ obeys bounds analogous to (60), namely
$g(v)\leq\mathcal{M}(v)\leq g(\zeta^{-1}(v))~.$ (111)
@@
$\int\text{d}^{d-2}y_{\perp}$ $\int\text{d}v\,\mathcal{M}(u,v)\langle T_{vv}(u_{0},v,\vec{y}_{\perp})\rangle\geq$
$-\frac{\beta_{d}c_{\scriptscriptstyle\mathrm{T}}V_{\perp}}{2\pi}\int\text{d}v\left(\frac{\partial_{v}\mathcal{M}(\zeta)\Theta(\partial_{v}\zeta-1)+\partial_{v}\mathcal{M}(v)\Theta(1-\partial_{v}\zeta)}{\epsilon_{u}^{\frac{d-2}{2}}(v-\zeta)^{\frac{d}{2}}}+O(\epsilon_{u})\right)$ (112)
where for brevity of notation we have left the $v$ dependence of $\epsilon_{u}$ and $\zeta$ implicit.
Again, much like our simplified 2d QEI, (61), (3.2) is now expressed in terms of a single smearing function, $\mathcal{M}$, and two functions, $\zeta$ and $\epsilon_{u}$ parameterizing choices of nearly null intervals. For a given $G$, which is fixed by the theory in question, it is unclear to us whether any generic smearing function $\mathcal{M}$ can be realized by a choice of $g$ appearing in (3.2). If so, then (61) and (3.2) would represent a very tantalizing result: a universal family of QNEIs for interacting CFTs which, up to an overall constant, are independent of the particular details of a CFT.
@@ 4 Discussion
In this paper, we have derived new quantum null energy inequalities, i.e. lower bounds on $\langle T_{vv}\rangle_{\psi}$, the null-null component of the stress tensor, integrated over a spacetime region in interacting quantum field theories. While our study of QNEIs is partially motivated by their potential applications to semi-classical gravity (e.g. as input in a semi-classical singularity theorem) such lower bounds on null energy are, of course, of intrinsic interest in the study of interacting QFTs. In particular, the foundations of our QNEIs are deep connections between energy and entropy: in our derivation, we utilize QNEC, strong subadditivity, and the vacuum modular Hamiltonians of null strips. More specifically, given a positive smearing function, $g$, we bound the stress tensor integrated against a smooth kernel, determined by the properties of the modular Hamiltonian of a null strip, in terms of a state-independent functional of $g$. In the case of $d>2$ dimensions, our bounds are the first such state-independent QEIs integrated over a finite null segment in a broad class of interacting QFTs, namely CFTs.
@@
where a precise form of $m$ can be found in (59). In contrast to QNEC, these lower bounds are state-independent, because the right-hand side does not depend on the state $\psi$. And, in contrast to ANEC, these bounds are localised, in the sense that we can choose $g$ such that we are integrating $\langle T_{vv}\rangle$ against a bump function on the left-hand side.
We have an infinite family of QNEIs, one for each choice of $\zeta$, which is both powerful but also increases the complexity of the inequality. The choice of $\zeta$ affects both the definition of $m$, the correction to the smearing function, and the value of the lower bound and arises from a choice of null intervals appearing in intermediate stages of our derivation. However, ultimately $\zeta$ is an invertible function obeying (48), but otherwise arbitrary. For a given $\langle T_{vv}\rangle$, there does exist a $\zeta$ that makes this intermediate stage inequality as tight as possible, but, for general $\langle T_{vv}\rangle$, there is no analytic solution for this $\zeta$ (see Sec. 2.2.4). We explicitly determined the QNEI for a few example choices of $\zeta$. It would be interesting to see whether the general form of our QNEIs can be simplified.
@@
$\begin{split}\int\text{d}^{d-2}y_{\perp}&\int\text{d}v\mathcal{M}(v)\langle T_{vv}(u_{0},v,y_{\perp})\rangle\\ &\geq-\frac{\beta\,c_{\scriptscriptstyle\mathrm{T}}}{2\pi}\int\text{d}^{d-2}y_{\perp}\int\text{d}v\left(\frac{\partial_{v}g(v)}{\epsilon_{u}(v)^{\frac{d-2}{2}}(v-\zeta(v))^{\frac{d}{2}}}+O(\epsilon_{u})\right)~,\end{split}$ (114)
@@
where $\mathcal{M}$ is defined in (3.2). As for $d=2$, this inequality is state-independent, [^15] To be precise, we have only shown state-independence for powers of $\epsilon_{u}$ between $\epsilon_{u}^{-\frac{d-2}{2}}$ and $\epsilon_{u}^{0}$. We expect that the subleading, positive-power orders in $\epsilon_{u}$ in the bound (114) are not state-independent, that they depend on the expectation values of the single and multi-copy operators with twist $\tau>d-2$ (greater than that of the single-copy stress tensor) that appear in the twist-twist OPE (3.2). and the integral over $\langle T_{vv}\rangle$ is localised (in the $(u,v)$ plane). Unlike for $d=2$, we integrate $\langle T_{vv}\rangle$ over the transverse directions $y_{\perp}$. Additionally, this bound depends on a second function, $\epsilon_{u}$, which parametrises how close we take our strip regions to being null in an intermediate step of the derivation (see Sec. (3.2)), and we have to work perturbatively in this function, to make use of the universal form of $\Delta K$ for null strips. To our knowledge, these are the first QEIs, integrated over a finite null region, proven for higher-dimensional interacting QFTs.
@@
While we view this work as an important first step in the study of QNEIs in interacting theories, there is still much more to explore:
@@
  • Our QNEIs in higher dimensions are localised in the $(u,v)$ plane, but not in the transverse space, so they are not truly localised bounds on the stress tensor. Additionally, this integration strictly trivialises the right-hand side of our bounds (albeit in a physically uninteresting way). It would obviously be interesting and useful to try to leverage the techniques we’ve established here to find bounds that are also localized in this transverse direction. At a technical level, it was necessary to have a flat entanglement cut in the transverse directions to make use of the results for $\Delta K$ for a null strip given in Bousso et al. (2015). Having the expression for $\Delta K$ for null strips with “wiggly edges” would allow us to utilize local variations and the QNEC in its local form, (67). However, this expression for the modular Hamiltonian is not known outside of free theories. The modular Hamiltonians of infinite null half-sheets with wiggly edges do admit closed form expressions Faulkner et al. (2016a); Casini et al. (2017) however we have so far not been able to leverage this expression to derive any new interesting bounds. Regardless, it is worth revisiting and reviewing this approach more carefully in the future. We expect that it is possible to localise the QNEI (114) in the transverse directions. Removing the integrals over $y_{\perp}$ in (114) gives a reasonable, transversally-localised QNEI conjecture. We leave the proof for future work.
@@
    It is also useful to look for alternative approaches for deriving a truly local QNEI. One potential approach is to try to realize the stress tensor in the OPE of two operators. This is essentially a more sophisticated form of the free field theory statement that $T_{vv}$ is the square of two operators. One obvious hang-up to this approach is that operator spectra in CFTs are not universal and so there is potential ambiguity in which OPE we try to realize stress tensor. To address this, we can once again utilize entanglement-based techniques. Namely, defect CFTs living along twist defects defining replica path-integrals admit a universal set of displacement operators with the stress tensor appearing in their OPE channel Balakrishnan et al. (2022). One might speculatively hope that a reflection-symmetric OPE of such displacement operators could provide a universal and local QEI in interacting CFTs.
@@
  • We have worked in Minkowski spacetime, both for simplicity and because our QNEIs are only proven for theories and background geometries for which our inputs are proven, QNEC in particular. The proof of QNEC given in Balakrishnan et al. (2019) is valid for all relativistic QFTs in Minkowski spacetime. However, there are also proofs of QNEC on curved backgrounds, for a smaller set of QFTs, such as in Bousso et al. (2014); Fu et al. (2017). It would be advantageous, particularly with the application to semi-classical gravity in mind, to generalise our work to curved backgrounds. One obstacle to overcome is the transverse integration of the previous bullet point. Because the integration region of our bound in $d>2$ dimensions could detect global structures of a background, we might incur large curvature corrections or even topological obstructions to our bound. A more local expression of the form mentioned in the previous bullet would allow us to work perturbatively when smearing over regions smaller than the typical curvature scale.
@@
  • Our derivations take QNEC as one of their inputs. In the same way that, while ANEC can be proven by integrating QNEC along a complete null line, it can also be proven directly using causality Hartman et al. (2017) and monotonicity of relative entropy Faulkner et al. (2016a), so we expect that our QNEIs can be proven without taking QNEC as input, perhaps using the causality and monotonicity of relative entropy.
@@
  • It would be interesting to extend our work to QFTs coupled to gravity. One intriguing possibility is that gravity itself provides a natural regulator of negative energy densities. This is essentially the core idea of the SNEC (4) mentioned in the introduction of this paper. However, as mentioned there, collective evidence for or even proving such QEIs is a hard task because they require coupling to gravity as input. Existing proofs of QNEC, such as Balakrishnan et al. (2019), assume a fixed background geometry, but the original statement of QNEC came as a corollary of the QFC Bousso et al. (2016b), and the motivation of that paper was to study QFTs coupled to gravity. It would be interesting to see whether QEIs for QFTs coupled to gravity can be proven directly from the QFC. We are actively working in this direction and will report results in a future publication Fliss and Rolph .